hi..pls.help me on my home work on mixture problem!......*How many gallons of water is added to 10 gallons of a 40% salt solution to make it a 15% salt solution?
Answer by Alan3354(69443)

These are all similar:
============================
How much decaffeinated coffee(assume to be 0% caffeine)and 40% caffeine coffee should be mixed to produce 20 mugs of 30% caffeine coffee?
------------
d = decaf
c = regular
----------
d + c = 20 (total mugs)
0*d + 40*c = (0 + 40)*30 (total caffeine)
40c = 1200
c = 30 mugs
d = 10 mugs